OpenStudy (anonymous):

Find the volume of the solid obtained by rotating the region bounded by y = 1/x^4, y = 0, x = 1, x = 5 about the y-axis

OpenStudy (anonymous):

@zepdrix I learned the disks/washers but not shells

OpenStudy (anonymous):

@zepdrix yes

OpenStudy (anonymous):

@zepdrix a little bit confusing, but i am haning on

OpenStudy (anonymous):

@zepdrix i got a huge decimal 4.8001006.......

OpenStudy (anonymous):

@zepdrix this is online homework, so i have an unlimited number of attempts to try to get the right answer and then it will tell me when i get it right

OpenStudy (phi):

FYI, Cylindrical shells seems to be a nicer approach  integrate from x= 1 to 5 \[ 2 \pi \int_1^5 x^{-3} \ dx = 2 \pi \cdot - \frac{1}{2} x^{-2} = -\frac{\pi}{x^2} \bigg|_1^5= \\ - \frac{\pi}{25} + \pi = \frac{24}{25} \pi = 0.96 \pi\] using discs, you should get the same answer, though it is a bit more complicated.